OpenStudy (anonymous):

Suppose f and g are prices of traded securities that depend on the same source of uncertainty. Their evolution is governed by the following geometric Brownian motions: df = αfdt + βfdz dg = γgdt + δgdz Let P be the relative price of g in terms of f: P == f/g (a) (10 points) Find the expression for dP [Hint: Proceed in the following steps. (1) Define x == In(P) and use the Lognormal Property (or Ito's Lemma if you prefer) to find dx. (2) Note that P = e^x: and use Ito's lemma to find dy = dP. (b) (5 points) What restriction on the parameters α, β, γ, and δ would make P a martingale? Explain.Suppose f and g are prices of traded securities that depend on the same source of uncertainty. Their evolution is governed by the following geometric Brownian motions: df = αfdt + βfdz dg = γgdt + δgdz Let P be the relative price of g in terms of f: P == f/g (a) (10 points) Find the expression for dP [Hint: Proceed in the following steps. (1) Define x == In(P) and use the Lognormal Property (or Ito's Lemma if you prefer) to find dx. (2) Note that P = e^x: and use Ito's lemma to find dy = dP. (b) (5 points) What restriction on the parameters α, β, γ, and δ would make P a martingale? Explain.

OpenStudy (jamesj):

x = ln(P) = ln(f/g) = ln f - ln g Then by Ito's lemma, dx = ( αf/f dt + βf/f dz) - (γg/g dt + δg/g dz) = (α - γ) dt + (β - δ) dz Now I'll let you figure out how to get back to dP using Ito's lemma again.

OpenStudy (anonymous):

JamesJ, could you please describe how to get back to dP by using Ito's lemma because I am stuck on that part. Please be detailed because I am so stuck on this little part. It would be most helpful so I would very much appreciate your effort!